The volume of a rectangle chineses box is given by the formula V=LWH. The volume of the box shown on the right is 396 cu units. Find the width and length. Height 4; width is X; length is X+2. Please help me solve this.
Answer by ankor@dixie-net.com(22740)

The volume of the box shown on the right is 396 cu units.
Find the width and length. Height 4; width is X; length is X+2.
:
x*(x+2)*4 = 396
divide both sides by 4, results:
x(x+2) = 99
:
x^2 + 2x - 99 = 0
factors to
(x + 11)(x - 9) = 0
the positive solution
x = 9 units is the width,
you can find the length
:
:
Check it by multiplying L*W*W